Brandom’s Philosophy of Logic

Michael Rescorla

Robert Brandom’s ”Making it Explicit” advances a novel conception of logic. According to Brandom, logical locutions allow speakers to analyze linguistic practice by making explicit the inferential norms which implicitly govern that practice. For instance, a speaker can signal acceptance of the inference from p to q by using the conditional, in particular by affirming the sentence ”If p, then q.” Of course, the classical material conditional is not well-suited for this role. But, for precisely that reason, Brandom denies that natural language ”if-then” corresponds to the classical material conditional. Brandom introduces a new kind of conditional, one more appropriate for his purposes. He then extends his analysis to other logical connectives, such as negation, identity, and the quantifiers. In my paper, I criticize Brandom’s conception of logic. In my view, logical locutions do not serve merely to facilitate analysis of linguistic practice. They serve a much more general and important purpose: they allow us to express complex views concerning any topic whatsoever. A language stripped of logical connectives would be far more expressively barren than Brandom seems to appreciate. I develop this line of criticism by reflecting upon the compositionality of linguistic meaning, a phenomenon which I argue that Brandom fails to explain satisfactorily. I further buttress my case by considering Brandom’s own paradigmatic example, the conditional. I argue that this example is far more problematic than Brandom acknowledges. In particular, the actual linguistic data undermine Brandom’s claim that conditional locutions serve primarily to facilitate discussion of inferential norms. I conclude with some general remarks about my discussion’s import for Brandom’s overall philosophical project.

Works Cited

Brandom, Robert. ”Making it Explicit.” Cambridge: Harvard University Press, 1998.

Address: Harvard University, 44 Lincoln Parkway, Apt. 1, Somerville, MA 02143, USA

E-mail: rescorla@fas.harvard.edu